Crew 211 Science Report 25Apr20189

[title Science Report – April 25th]

[category science-report]

Science Report 25 April 2019
Crew 211 – UCL to Mars
Submitted by Science Officer Maxime Bernard

1) Commander Carl Dahlqvist

Working on the software to allow the main module to receive data from the secondary modules.

2) XO Officer

Coding a trilateration algorithm to compute the position of a target from the respective distances from that target to the 3 antennas.

3) HSO Officer Benjamin Flasse

Every morning and every evening, the HSO still gather data about the mensurations, weight, brain awakeness, tension and pulse of the crew.

4) Crew Astronomer Eleonore Lieffrig

Working on the images she got from the remote telescope.
In the meantime she started to work on the Brownian motion operator.

5) Crew Engineer Julien Amalaberque

No major milestone reach for now.

6) Crew Biologist Chloé Peduzzi

She checked the spirulina through the microscope and confirmed it was the right species (Arthrospira Platensis). Moreover, she observed their spiral shape which indicate their healthy state.
She then proceeded to the aeration of all the spirulina culture.

7) Crew Journalist Nathan Pechon

We made a meeting during which he presented us law articles about birth and immigration rights on a martian colony and we had to vote these articles.

8) Crew Scientist Maxime Bernard

He assembled his detector and started the acquisition of data from muon interactions

Science Report – April 24th

Science Report 24 April 2019 Crew 211 UCL to Mars Submitted by Science Officer Maxime Bernard 1) Commander Carl-Henrik Dahlqvist The 8 CubeSat cases are assembled and two of them are already functional. He is working on the code to increase the size of the cubelanders network. 2) XO Officer All four antennas are assembled and fixed to the poles. At the moment, he is still working on the software side of the project. 3) HSO Officer Benjamin Flasse Every morning and every evening, the HSO still gather data about the mensurations, weight, brain awakeness, tension and pulse of the crew. In the meantime he is also making an inventory of all the first aid kits to ensure they are all complete and reliable. 4) Crew Astronomer Eleonore Lieffrig Our astronomer managed to take pictures of the sun with the telescope and solar eruptions could be seen. She is also preparing the observation of a quasar using the remote controlled telescope. 5) Crew Engineer Julien Amalaberque No major milestone reach for now. 6) Crew Biologist Chloé Peduzzi The crew engineer and the XO managed to inverse the transformer to allow Chloé to use her LED light for her first sample of spirulina. She also placed a temperature sensor near the spirulina to make sure the temperature will never exceed 30°C. She then started her second experiment in which she is going to plant tomato and bean seeds in different medium composed of different quantities of compost/martian soil with mycorrhiza and/or hydrogel which are going to be dry or not. It will finally constitute a total of 16 different medium. 7) Crew Journalist Nathan Pechon We made a meeting during which he presented the voting rules and the different possible regimes available according to the talks we all had with him earlier. During this meeting we voted for the political regime we thought would suit the best an early martian colony. 9) Crew Scientist Maxime Bernard The detector finally arrived in Hanksville and Maxime together with Shannon picked it up around 6:30 pm.

	

Science Report – April 22nd

Crew 211 – UCL to Mars

Submitted by Science Officer Maxime Bernard

Commander Carl-Henrik Dahlqvist

Our commandant started to assemble the structure of his CubeSat that he printed prior to his arrival on Mars.

XO Officer

Simon started to build his wood stands on which he is going to fix his antennas.

HSO Officer Benjamin Flasse

This morning Benjamin gathered the data from the sleep monitoring he did on Eleonore.

Right after that, he measured everybody’s weight, tension and brain awakeness.

In addition to that, he took measures like fat density,bone quality and water repartition intra/extra cellular using Biody Xpert to monitor our body evolution along our stay here.

He spent the rest of the day making the inventory of what was available in the first aid kits. Most of the compresses were outdated and others were opened. The purchase of new medical supplies should be prioritized and more info will be available in one of the upcoming HSO reports.

Crew Astronomer Eleonore Lieffrig

Our astronomer was able to use the solar telescope for the first time today in manual mode.

Crew Engineer Julien Amalaberque

He is checking the parameters of an Intel RealSense D435i depth camera in order to evaluate which amount of stream data the positioning algorithm will be able to handle without being overwhelmed.

Crew Biologist Chloé Peduzzi

She spent the morning one the science dome preparing her culture medium in which she adds strains of spirulina.

She then disposed them in the greenhouse under two different light conditions before realizing it was way too hot in there. Therefore, she instead disposed them into the Science Dome. She encountered another problem we she realized she was sold the wrong transformer, 240V to 120V instead of 120V to 240V.

She went on EVA in the afternoon with three other crew members to gather soil samples.

Crew Journalist Nathan Pechon

Today, Nathan started to ask us questions about what we thought would be a good juridical baseline of a martian colony.

He then worked on the data he gathered during our talks to work on his first law article about what type of political regime should be adopted on Mars

Crew Scientist Maxime Bernard

Still waiting for the delivery of his muon detector that should have arrived on the 19th of April… The parcel seems to be somewhere between Salt Lake City and Hanksville.

Science Report – April 18th

Science Report 18 April 2019
Crew 210
Submitted by Paul Sokoloff

Today we continued to process the soil samples taken at each of the 12 transects we conducted over the last few days. In the Science Dome, we measured the pH and salinity and characterized the texture of the soil samples. These data will be used to characterize the ecotones of the deserts around MDRS in an upcoming paper.

We continued to collect new plant specimens opportunistically, including an annual invasive grass, (Annual Wheatgrass – Eremopyrum triticeum) newly recorded for MDRS.

We plan on continuing our lab work and wrapping up data recording for the plant project tomorrow.

Science Report – April 17th

Continuing our vegetation ecology and biodiversity project, we plotted transects and surveyed quadrats at two new sights today: Copernicus Valley and Hab Ridge. As yesterday, we surveyed the vegetation along three transects and fifteen quadrats per site, recording abiotic characteristics and the vegetation of each site.

Additionally, we continued to collect voucher specimens for new vascular plant species encountered at each site, adding 15 new collections and ~5 new species for the MDRS operational area.

This evening we plan on entering data and processing data stemming from these surveys.

Science Report – April 16th

This afternoon crew 210 began to conduct vegetation quadrat surveys as a part of our local plant and lichen biodiversity survey. At two sites pre-selected from the MDRS topographic map (both east of the hab along Cactus Road) we plotted three 10-meter transects, and surveyed five 1-meter-square quadrats along each transect in a zig-zag pattern.

For each transect we recorded GPS coordinates, air and soil
temperature at the center of the transect, measured the height of the tallest plant on the transect, and took a central sample for soil moisture and pH (these latter variables will be measured in the lab tomorrow).

Within each of the five quadrats, we estimated percent cover of vascular plants, and named the top three species (by abundance) inside the square-meter. Any new plant species encountered were collected for identification and to add to our plant biodiversity data.

Tomorrow we will finish the quadrat surveys at three additional pre-determined sites, north and west of the hab, and will continue our plant and lichen collection efforts.

Regards,
Paul, on behalf of Crew 210

Science Report – April 14th

Sol 2 – Science Summary

Crew 210 sampled vascular plants and lichens from three sites across the MDRS operational area today; the Valley of the Stars, Salt Creek, and the vicinity of the Burpee Dinosaur Quarry. We collected between 15-20 samples, many of which were species not recorded during previous floristic work at the station in 2014. We will press the samples later this evening to preserve them for later identification. Tomorrow, if time permits, we will continue our sampling in the area just around the hab.

Science Report – March 23rd

Crew 207

Research Report

Submitted by Crew Scientist Tiffany Ni

3D Printing Projects

Today is the last full day of MDRS for myself and our crew HSO. Our 3D-printing experiments continue to progress. We have completed printing all of our labware, International Space Station medical inventory items, STEM teaching aids, and are now printing some things here and there around MDRS, as well as other miscellaneous 3D printed projects.

To date, we have printed:

1. Western blot combs (18 well combs for making gels)
2. Microscope slide holder (reusable holder for square coverslips)
3. Sieve mesh (2.0mm holes)
4. Sieve structure (for holding the mesh net)
5. Mobile phone microscope clip
6. Sample box (for 96 tubes)
7. Stackable freezer bin
8. Salt spray chamber drain
9. Triangular straws (in a rainbow of colors)
10. Sieve mesh (5.0mm holes)
11. Sieve structure (for holding the mesh net)
12. Slide box holder (for storing microscope glass slides)
13. Test tube rack
14. Test tube holder
15. Sample rotator components
16. Small test tube holder (for the freezer)
17. Heat gun nozzle mount
18. Tongue depressors (in a rainbow of colors)
19. Stoma bag clip (used as a chip bag clip)
20. Full set of fraction tools (mathematics educational tools)
21. Funnels (small, medium, large)
22. Construction blocks
23. 12 tube separator (for Eppendorf tubes)
24. The side part of thoracentesis trainer
25. Customized finger splint (originally printed on the ISS0

During our time at MDRS, our 3D printed straws proved incredibly useful for our crew members. It has helped greatly in our attempts to reduce our water consumption. At MDRS, we have also demonstrated that a number of our 3D printed labware items are functional – as validated by two basic science research students.

Even though we will be leaving tomorrow morning, we will continue to stay in contact with our MDRS crew members and send 3D printable files for incoming crew members to print! (Perhaps some reusable bubble tea straws?)

Signing off,
Tiffany Ni
Crew 207 Scientist

Science Report – Mar 06th

Crew 206 Science Report 06-Mar-2019

You’re lucky, today we wrote our first science report of the mission. You’ll have the chance to see our first results !

The attached piece is the graphs of our first week on water monitoring.

Regards,

Crew 206
Submitted by HSO/Science Officer Cerise Cuny

1. Microbiological water analysis (Aquapad) : I have done 3 series of 3/4 1 mL aquapads since the beginning of the mission. The number of bacteriological colonies is higher than the sanitary criteria (100 UFC) but lower than the infectious limit (300 UFC). As the water we extract in Hanksville is without a doubt adapted to drinking, maybe the bacterias proliferate in the tank as it is exposed to the sun all day.

Serie 1 : 117 UFC in average after 48h of incubation
Serie 2 : 227,7 UFC in average after 48h of incubation
Serie 3 : 192 UFC in average after 48h of incubation

2. Water Monitoring : For now, our way of monitoring our water consomption is way more efficient than last year Supaero Crew’s (Crew 189). Nevertheless the system can still be improved. We have a new method to do the dishes that is more economical. See attached PDF for numbers.

3. Simulation space suit : It was brought to MDRS to be compared to the actual suits and the conditions of the simulation. They were able to put into light the imperfections of their suit but also its advantages and they oriented the future development process.

4. EVA tracker app : We are still facing issues.

5. Music for plants : The biggest issue is an equipment one : the UV lamp the Greenhab Officer brought didn’t work due to the differences in US and EU electrical system. He used the UV lamp that was in the science Dome which is not powerful enough. The experiment may not work as planned.

6. The human factors experiments won’t be described here in order to avoid bias.

Water monitoring.pdf

Science Report – February 17th

Crew 205 – International Emerging Space Leaders

Submitted and prepared by GreenHab Officer Nathan Hadland

1.) ISRU Study: Continued characterization of the regolith samples have produced interesting properties. The pH of most of the samples have yielded a pH that is neutral or slightly alkaline with the exception of one, which is slightly acidic. These values suggest that the substrates may be good candidates for plant growth and retaining nutrients. Higher pH substrates cause the Hoagland’s #2 hydroponic nutrient supplement to precipitate out. The probes in the Science Dome to investigate conductivity and redox potential are not operational, so this property will not be investigated until returning to Florida Tech. Using an agar media plug, germinated Arabidopsis thaliana seeds were transferred from the agar germination plate into regolith samples. However, the regolith is exhibiting interesting wettability properties, largely drying out within several hours despite ample watering, effectively killing the fragile Arabidopsis seedlings quickly overnight. The Arabidopsis seeds were planted again this afternoon, with more plants per sample and care to arrange the regolith around the plug to prevent quick dehydration. Dead plant material from the GreenHab was taken and dried using an oven in the Science Dome and subsequently weighed for the cellulosic ethanol ISRU study. We will begin a study soon mixing the regolith samples with potting soil to try to improve the success of this study. See attached pictures to see experimental design.

2.) GreenHab: The inventory of equipment in the Science Dome and GreenHab as well as the plants currently growing is nearing completion and we have begun a list of the seed bank. We suggest these lists to be available to crews prior to rotation and checked and updated while at MDRS. This will improve the ability of crews to design experiments and mission plans effectively and improve the efficacy and safety of all participants involved. Additionally, the sugar snap peas appear to be stagnating, so a nutrient supplementation solution (Miracle Grow) was diluted using the industry standard and applied to the base of the peas along with a few other species sparingly.

3.) EVA Project: The success of an extravehicular activity (EVA) depends highly on both the EVA crew and CapCom crew being aware of the planned objectives and the intended route. This means that both parties should conduct the planning of the EVA together, in order to understand what is to be accomplished and what hazards could change the EVA’s outcome. Crew and CapCom should also be in contact during the EVA for both safety and additional information on EVA progress. The following post-EVA checklist will allow crews of planetary exploration analog sites going on EVAs to expose deficiencies in both the planning and execution of the EVA. This process allows planners and team members to improve their process of preparing for an EVA. This will lead to more effective and safe EVAs.

4.) SOP and Guidelines for Future Crews: So far, the IESL Crew (Crew 205) have drafted Standard Operating Procedures (SOPs) that provide checklists for leaders, HSOs and support crew for before and after EVA missions. These SOPs will improve efficiency, reduce ambiguity and streamline the processes involved in preparing for and executing surface operations safely and effectively. Some of the overwhelming themes contained in these SOPs are:

o Ensuring that requisite special mission equipment is secured and checked prior to EVAs

o Ensuring that participants understand both the objectives and scheme of movement for the EVA

o Ensuring that durable equipment (EVA suits, tools, etc.) is operable and mission-ready prior to EVA, and is properly recovered afterwards, with maintenance needs immediately annotated

o Ensuring that potential risks have been identified and mitigated prior to departure

o Ensuring that all of these requisite tasks are completed by every EVA team of every crew as an operational standard, not a mere suggestion

5.) Astronomy: One week has passed since the start of crew 205 rotation. Two observations were submitted, but the weather prohibited the completion of one observation. The target observed was SY MON, while we called the other target MDRS Target 1. 4 photometry measurements were performed for SY MON as per the below table. The Magnitude Value of 14.132 was submitted for the AAVSO website.

Mag Value Comp Stars Labels Check Star Label Error Comment
14.147 93 111 122 134 148 No Check Stars were used 0.01568 This measurement was repeated because of no check star was used, CCD setting was wrong and there are better ways to select comp stars.
13.918 140 143 132 97 Check star used indicates bad measurement or check star itself was saturated Measurement repeated
14.118 140 143 132 No check star because this is manual calculation It was found that Comp 97 positioned outside the straight line of the curve, thus it was decided to exclude it.
14.132 140 143 132 122 0.0185 Data was submitted to AAVSO